how do u do 90-(3x-7)=11x-1
Found 2 solutions by rfer, Jk22:
Answer by rfer(16322)

90-3x+7=11x-1
98=14x
x=7

Answer by Jk22(389)

90-(3x-7)=11x-1
97-3x=11x-1 | +1+3x on both side
98=14x | div by 14, 7*14=70+28=98
7=x